find three consecutive numbers such that the sum of the second and the third numbers exceeds the first by 14.
Answer by mananth(16949)

let the three consequtive numbers be x, x+1, x+2

x+1+x+2=x+14
2x+3=x+14
-x
2x-x+3=14
x+3=14
-3
x=11
11,12,13